# Ferraz, Cappart, and Vidal

### v1

@@ 6.3 Results – Prediction Accuracy
Figure 3: Visualization of the ground truth and estimated district costs on a subset of districts for Bristol and London considering central depot.
Finally, Figure 3 provides graphical representations of the ground-truth costs computed by SAA, along with their confidence intervals and the predicted cost values from the estimation methods. The light yellow area represents the cost provided by SAA and its 95% confidence interval (due to the finite number of samples in the test set). The darker continuous curve corresponds to SNN, and the lighter continuous curve corresponds to the GNN. To facilitate the visualization, we only display the results for a random subset of 300 districts ordered by increasing ground-truth cost. The figure displays these graphs for two different metropolitan areas (Bristol and London), considering centered depots and three possible configurations for the $n$ and $t$ factors. Generally, we can observe that GNN provides better estimates than SNN for each scenario. Beyond this, comparing the figures with $t=3$ and $t=12$, we notice that lower $t$ values lead to much smaller errors, i.e., GNN and SNN estimates are much closer to the true cost. Larger values of $n$ also tend to diminish prediction accuracy, but this effect is less marked than when changing the district sizes.
6.4 Results – Operational Efficiency of the Districting-and-Routing Solutions
@@
This section evaluates the influence of various cost estimation methods within the context of a districting-and-routing solution approach. We conduct our experiments with four estimation methods: BHHD, FIG, SNN, and GNN. The iterated local search metaheuristic (see Section 5) has been designed to be generic and accommodate any cost estimation algorithm. This versatility enables us to switch estimation approaches and gauge their effects on the districting-and-routing solutions. We refer to the variations of the ILS employing different cost estimators as GNN-ILS, BHHD-ILS, FIG-ILS, and SNN-ILS. We systematically evaluate the district quality using SAA in conjunction with LKH across 500 test scenarios to obtain cost estimates. All methods have the same time budget available: 180, 600, and 1200 seconds for instances of 60, 90, and 120 BUs, respectively.
@@
Comparisons with full-knowledge solutions on smaller instances. For the smallest instances, it is possible to enumerate all the possible districts, evaluate their costs using the SAA approach on the test scenarios, and solve the graph partitioning formulation of Equations (36). This approach gives us full-knowledge solutions that can be used as a bound on the best possible cost ($z_{\textsc{saa-test}}$). However, this requires a massive computational effort. As an illustrative example, let us consider an instance with $n=60$ and $t=6$. On average, approximately 700,000 feasible connected districts exist that adhere to the constraints of the problem. Consequently, $500\times 700,000$ TSPs must be solved (one for each potential district and test scenario) to estimate ground-truth costs and subsequently solve the related set-partitioning formulation. Given that each TSP solution requires roughly 0.1 seconds, the total computational effort for the full-knowledge optimal approach is roughly 405 CPU-core days for each instance. We could undertake this computation only for instances with $(n,t)\in\{(60,3),(60,6),(90,3),(90,6),(120,3)\}$ using a large computer grid. We highlight that any search method relying on SAA encounters similar limitations since any solution-cost evaluation is a time-intensive process. Relying on a supervised learning approach to estimate the costs alleviates this issue.
@@
General comparison. Considering all the instances, full-knowledge solutions are in most of the cases unavailable, and therefore we present the percentage solution-cost gaps for each method (BHHD-ILS, FIG-ILS, and SNN-ILS) in comparison to GNN-ILS, employing out-of-sample evaluation on the test set. Table 6 details these gaps across all instances.
As seen in this experiment, the GNN-ILS consistently demonstrates superior performance compared to other methods, often by a significant margin. The average gap ranges between $10.12\%$ and $13.57\%$ depending on the method and depot configuration. This underscores the substantial impact of accurate district cost estimations on the performance of strategic districting-and-routing solutions. Additionally, we observe that our approach’s relative performance strengthens as district size increases. This result underscores the ability of the graph neural network to exploit the inherent structural characteristics of the districting problem, resulting in improved estimations for larger districts.
@@
Visual analysis. Figure 4 compares the solutions of GNN-ILS and SNN-ILS on two instances based on the Manchester metropolitan area, the first case (Manchester_C_120_6) with a target district size of $6$ and a centered depot, and the second case (Manchester_SE_120_20) with a target district size of $20$ and an off-centered depot in the south-eastern part of the area. Each district is distinguished by a different color, and the depot is represented with a dot. In those two cases, the operational cost of the GNN-ILS solution is significantly inferior to that of the SNN-ILS solution, with improvements of 10.08% and 11.57%, respectively. The solutions of the GNN-ILS are also (i) more compact and (ii) exhibit a geometry in which the districts are elongated towards the depot location. As discussed further in this section, this latter characteristic is intuitive and critical to achieve good performance. None of the features of BHHD, FIG, or SNN measure this geometrical characteristic; therefore, these structures do not naturally emerge. In contrast, the GNN learns from past good combinations of BUs and has the ability to learn to promote such characteristics.
Refer to caption (a) SNN-ILS solution on Manchester_C_120_6 Compactness Score: $0.327$, Cost: $794.15$
Refer to caption (b) GNN-ILS solution on Manchester_C_120_6 Compactness Score: $0.392$, Cost: $714.18$
Refer to caption (c)SNN-ILS solution on Manchester_SE_120_20 Compactness Score: $0.245$, Cost: $511.19$
Refer to caption (d)  GNN-ILS solution on Manchester_SE_120_20 Compactness Score: $0.360$, Cost: $452.07$
@@
Geometrical characteristics of good districts. Compactness scores have often been used as a proxy measure to maximize the efficiency of routing operations within districts (Benzarti, Sahin, and Dallery 2013, Bruno et al. 2021), and Figure 4 illustrated that cost-efficient districts are generally compact. However, a good operational performance goes beyond simple compactness measures, and therefore is not fully aligned with this objective. Notably, Table 8 reports the compactness of the solutions generated using the ILS with the different estimators.
@@
[…]
@@
As seen in this table, when the depot is centrally located, the most efficient solutions in terms of cost (produced by the GNN-ILS) also feature the highest level of compactness, so it appears that compactness and routing efficiency are well aligned in that case. However, this holds to a lesser extent for large instances and when the depot is off-centered, as optimal vehicle routing problem solutions in such cases exhibit routes that are not circular but elongated towards the depot (e.g., see the solutions of the instances from Uchoa et al. (2017) accessible at http://vrp.galgos.inf.puc-rio.br/index.php/en/). Since such routing solutions are efficient, one can also expect to observe strategic districts possess similar geographical characteristics. These characteristics are visible in the solutions found with the GNN (see, e.g., Figure 4d). Interestingly, the GNN can guide the optimization towards such geometries and generally compact solutions without specific features measuring these characteristics. This highlights the flexibility of this supervised learning approach; as it gauges solution quality through imitation of known training districts, it can adapt to different solution characteristics and geographical elements without modifications or manual input. Therefore, this approach could be transferred to different settings (e.g., measuring delivery costs according to a different subproblem or considering additional constraints or network characteristics) by simply training on examples generated with a different cost.
@@ 7 Conclusions
This paper has introduced a novel approach for strategic districting-and-routing, relying on a GNN for district-cost estimation within an ILS aiming to find a good partition of a metropolitan area into districts. The GNN is trained through a supervised process, using pre-computed TSP costs for various districting configurations derived from the Lin–Kernighan heuristic. Our experimental analyses demonstrate that this approach generates efficient districts, leading to substantial economic gains of 10.12% on average (and exceeding 20% in some cases) compared to other commonly used methods for routing cost estimations. They also show that compactness is an insufficient measure of district effectiveness and that other geometric characteristics (e.g., elongated district shapes towards the depot) play a key role in determining long-term operational costs.
This research opens many promising research avenues. A natural direction for future work involves analyzing the applicability of GNN-based cost estimations for other stochastic problems that require time-intensive second-stage evaluations. One main advantage of the proposed GNN approach is its flexibility. It is driven by data (geographical features of the BUs and training examples) and does not require distributional information over customer requests. Consequently, it is fairly generic, and adapting it to other district-cost definitions (e.g., considering a different subproblem than a TSP, introducing additional constraints or network characteristics) boils down to replacing the approach used to generate labeled districts. Other research perspectives could focus on improving the learning methodology. The GNN has been trained to minimize prediction error, regardless of the downstream use of these predictions in the optimization problem. Performance could be enhanced with integrated learning, following the predict-then-optimize principle (Elmachtoub and Grigas 2022), or using end-to-end pipelines such as the ones pioneered in Dalle et al. (2022). Finally, to cut down training computational effort, one could consider generating examples and training the model on smaller datasets and evaluate its generalization to larger data at inference time, as done in Baty et al. (2023).

### v2

@@ 6.3 Analysis – Prediction Accuracy
Figure 3: Visualization of the ground truth and estimated district costs on a subset of districts for Bristol and London (with a central depot).
Impact of district geometry on estimation performance. This experiment evaluates how district geometry affects the performance of prediction models. We focus on two geometrical characteristics: compactness (defined earlier) and elongation toward the depot (defined below). Let $\sigma^{\texttt{N}}_{d}$ and $\sigma^{\texttt{O}}_{d}$ represent the standard deviations of customer positions in district $d$, measured along the axis between the depot and the district barycenter, and the perpendicular axis, respectively. The ratio $\sigma^{\texttt{O}}_{d}/\sigma^{\texttt{N}}_{d}$ indicates the district’s elongation: a high ratio reflects elongation perpendicular to the depot, while a low ratio suggests elongation toward the depot. To avoid districts that are near or include the depot, we retain only the 50% most distant districts from the test set. These districts are then split into four categories based on their compactness (or elongation) percentile value $v$ ($[v<1]$, $1\leq v<50]$, $50\leq v<99]$, $v\geq 99]$). For each method, we report the relative cost estimation error of districts within each category, as shown in Figure 4 for compactness and Figure 5 for elongation, using violin plots.
Figure 4: Performance analysis of the different estimators for districts with varying compactness
Figure 5: Performance analysis of the different estimators for districts with varying elongation
As shown in this experiment, the simpler baselines (BHHD, FIG, SNN) tend to underestimate costs in less compact districts and overestimate them in more compact ones. As a result, they often favor solutions with overly compact districts, although, as noted in Section 6.4, this may not always be optimal for districting and routing. A similar trend is observed with the elongation metric: districts elongated toward the depot (from where the delivery vehicles originate) are overestimated, while those elongated perpendicularly are underestimated. In contrast, the GNN produces consistent prediction errors across all metrics and categories.
Ablation analysis of the GNN features. We selected BU features commonly used in previous methods to train the GNN. Features like area and population were included both directly and as square roots, as seen in certain approximation formulas (e.g., BHHD). Although not exhaustive or minimalist, this approach leverages the neural network’s ability to identify and reinforce useful features, reducing the need for extensive feature engineering. To provide further insight into the importance of each feature, Table 5 summarizes the results of an ablation analysis, where the GNN was trained and tested with different subsets of the features. In this analysis, we retained the membership features defining the district for the prediction task, while selectively adding or removing one feature.
Table 5: Ablation Analysis: Prediction error when removing or adding GNN features
RMSE with all features: 3.14 RMSE with membership only: 4.67
- population 3.19 ++ population 4.06
- sqrt-population 3.24 ++ sqrt-population 4.11
- area 3.27 ++ area 3.73
- sqrt-area 3.21 ++ sqrt-area 3.60
- density 3.27 ++ density 3.74
- perimeter 3.22 ++ perimeter 3.57
- distance to the depot 3.24 ++ distance to the depot 3.94
- pop & sqrt-pop 3.28
- area & sqrt-area 3.29
Since the GNN is designed for training and application in a specific city, the membership variables alone enable the model to extrapolate known district costs from the training set to unseen districts, resulting in an average RMSE of 4.67. However, incorporating additional features significantly enhances this interpolation, reducing the prediction error to an RMSE of 3.14 when all features are used. As shown in the table, the individual impact of each feature is not immediately noticeable when deactivated, as the remaining features often compensate for the missing data. Still, when each feature is added to the membership information, it provides valuable insights that the GNN can leverage. The perimeter and square root of the area contribute the most to improving accuracy, reducing RMSE from 4.67 to 3.57 and 3.60, respectively. The importance of the square root of the area is expected, as it is a key parameter in approximation formulas like BHHD.
6.4 Analysis – Operational Efficiency of the Districting-and-Routing Solutions
@@
This section evaluates the influence of various cost estimation methods within the context of a districting-and-routing solution approach. We conduct our experiments with four estimation methods: BHHD, FIG, SNN, and GNN. The iterated local search metaheuristic (see Section 5) has been designed to be generic and accommodate any cost-estimation algorithm. This versatility enables us to switch estimation approaches and gauge their effects on the districting-and-routing solutions. We refer to the variations of the ILS employing different cost estimators as GNN-ILS, BHHD-ILS, FIG-ILS, and SNN-ILS. We systematically evaluate the district quality using SAA in conjunction with LKH across 500 test scenarios to obtain cost estimates. All methods are run on a single CPU thread and have the same time budget available: 180, 600, and 1200 seconds for instances of 60, 90, and 120 BUs, respectively. We note that this moderate time budget, up to 20 minutes in the largest case, is sufficient for convergence of the ILS (with optimality gaps below $0.1\%$) for each cost oracle, as shown in our calibration and validation experiments in the Appendix. Consequently, extending the runs yields no tangible benefit when fast district-cost predictors are used. Besides this, maintaining short run times is desirable in practice, as it enables rapid “what-if” analyses (e.g., considering different district sizes) and iterative decision support (e.g., reviewing, adjusting, or partially fixing districts) with a human in the loop, without retraining or long computation delays.
@@
Comparisons with full-knowledge solutions on smaller instances. For the smallest instances, it is possible to enumerate all the possible districts, evaluate their costs using the SAA approach on the test scenarios, and solve the graph partitioning formulation of Equations (36). This approach gives us full-knowledge solutions that can be used as a bound on the best possible cost ($z_{\textsc{saa-test}}$). However, this requires a massive computational effort. As an illustrative example, let us consider an instance with $n=60$ and $t=6$. On average, approximately 700,000 feasible connected districts satisfy the problem’s constraints. Consequently, $500\times 700,000$ TSPs must be solved (one for each potential district and test scenario) to estimate ground-truth costs and subsequently solve the related set-partitioning formulation. Given that each TSP solution requires roughly 0.1 seconds, the total computational effort for the full-knowledge optimal approach is roughly 405 CPU-core days for each instance. We could undertake this computation only for instances with $(n,t)\in\{(60,3),(60,6),(90,3),(90,6),(120,3)\}$ using a large computer grid. We highlight that any search method relying on SAA encounters similar limitations, as solution-cost evaluation is a time-intensive process. Relying on a supervised learning approach to estimate the costs alleviates this issue.
@@
General comparison. Considering all the instances, full-knowledge solutions are, in most cases, unavailable, and therefore we present the percentage solution-cost gaps for each method (BHHD-ILS, FIG-ILS, and SNN-ILS) in comparison to GNN-ILS, employing out-of-sample evaluation on the test set. Table 7 details these gaps across all instances.
As seen in this experiment, the GNN-ILS consistently demonstrates superior performance compared to other methods, often by a significant margin. The average gap ranges between $10.12\%$ and $13.57\%$ depending on the method and depot configuration. Interestingly, this performance gap is significantly larger than the differences in predictive errors (MAPE measured in Section 6.3), suggesting that optimization tends to amplify estimator errors by disproportionately selecting districts whose costs are underestimated. Additionally, we observe that our approach’s relative performance strengthens as district size increases. This result underscores the ability of the GNN to exploit the inherent structural characteristics of the districting problem, resulting in improved estimations for larger districts.
@@
Visual analysis. Figure 6 compares the solutions of GNN-ILS and SNN-ILS on two instances based on the Manchester metropolitan area, the first case with $20$ districts and a centered depot, and the second case with $6$ districts and an off-centered depot in the south-eastern part of the area. Each district is represented by a distinct color, and the depot by a dot. In both cases, the operational cost of the GNN-ILS solution is significantly inferior to that of the SNN-ILS solution, with improvements of 10.08% and 11.57%, respectively. The solutions of the GNN-ILS are also (i) more compact and (ii) exhibit a geometry in which the districts are elongated towards the depot location. As discussed in the next subsection, this latter characteristic is intuitive and critical to achieving good performance. None of the features of BHHD, FIG, or SNN measure this geometrical characteristic; therefore, these structures do not naturally emerge. In contrast, the GNN learns from past good combinations of BUs and has the ability to learn to promote such characteristics.
(a) SNN-ILS solution for Manchester
Centered Depot, $n=120$, $k=20$
Compactness Score: $0.327$, Cost: $794.15$
(b) GNN-ILS solution for Manchester
Centered Depot, $n=120$, $k=20$
Compactness Score: $0.392$, Cost: $714.18$
(c) SNN-ILS solution for Manchester
SE Depot, $n=120$, $k=6$
Compactness Score: $0.245$, Cost: $511.19$
(d) GNN-ILS solution for Manchester
SE Depot, $n=120$, $k=6$
Compactness Score: $0.360$, Cost: $452.07$
@@
6.5 Discussions
Geometrical characteristics of good districts. Compactness scores are often used as a proxy for routing efficiency within districts (Benzarti, Sahin, and Dallery 2013, Bruno et al. 2021), and Figure 6 illustrates that cost-efficient districts are generally compact. However, our results also show that operational efficiency extends beyond simple compactness measures and, therefore, is not always fully aligned with this objective.
In particular, Table 9 reports the average compactness of the solutions generated using the ILS with the different estimators. When the depot is centrally located, the most efficient solutions in terms of cost (produced by the GNN-ILS) also feature the highest level of compactness, suggesting that compactness and routing efficiency coincide well in this case. However, this holds to a lesser extent for large instances and when the depot is off-centered, as optimal vehicle routing solutions typically exhibit routes that are elongated toward the depot rather than circular (e.g., see the solutions of the instances from Uchoa et al. (2017) accessible at http://vrp.galgos.inf.puc-rio.br/index.php/en/). Because these elongated patterns are intrinsic to efficient routes, one can expect strategic districts to exhibit similar shapes.
@@ 6.5 Discussions
Off-Centered Depot $\{\texttt{NE},\texttt{NW},\texttt{SE},\texttt{SW}\}$ 3 0.367 0.369 0.369 0.368
@@
These characteristics are clearly visible in the solutions obtained with the GNN (e.g., Figure 6d). Interestingly, the GNN can guide the optimization towards such geometries, finding an efficient middle ground between compactness and elongation towards the depot, even without explicitly defined features measuring these properties. This underscores the flexibility of the supervised learning approach: as it gauges solution quality solely by imitation of known training districts. Consequently, the same methodology can be readily applied in various contexts, such as different routing subproblems, additional operational constraints, or network-based distance metrics, by training on district examples labeled with a different cost without any other architectural modification.
Experiments considering network-based distances. To further illustrate the flexibility of our approach, we evaluated its performance when distances are computed on the road network, therefore accounting for other real-world complexities such as tunnels, bottlenecks, and geographical barriers. Only minimal adaptation was required for this setting: the sole modification was to use road-network distances in the TSPs when generating the training districts. For this purpose, all-pairs shortest paths between delivery locations were computed using the contraction hierarchies algorithm of Geisberger et al. (2012) and Dibbelt, Strasser, and Wagner (2016), implemented in the RoutingKit library (https://github.com/RoutingKit). All the district-cost estimation models (GNN, BHHD, GIF, and SNN) were subsequently retrained or recalibrated on these new data to evaluate both their predictive accuracy and their impact on districting-and-routing performance. For brevity, these experiments were conducted on the subset of instances with $n=120$ basic units and a centrally located depot. The corresponding results are summarized in Table 10.
Table 10: Accuracy and operational performance of district-cost estimators under road-network distances
$\bm{n}$ $\bm{t}$ $\phi_{\mathrm{SAA}}^{\text{test}}$ RMSE MAPE Gap(%) vs GNN-ILS
GNN BHHD FIG SNN GNN BHHD FIG SNN BHHD FIG SNN
120 3 98.08 6.03 19.48 16.08 15.68 4.88 15.15 13.35 12.90 2.76 2.52 2.36
6 101.87 5.19 18.62 14.17 13.86 3.59 14.38 11.50 11.06 6.17 5.80 5.54
12 145.25 20.12 43.20 37.28 36.98 6.77 24.92 22.33 21.91 9.62 9.02 8.66
20 158.26 18.87 37.31 30.96 31.27 6.38 20.08 16.57 16.56 8.26 10.34 6.74
30 172.06 12.65 30.96 25.26 25.37 4.90 16.01 12.71 12.91 14.81 18.22 10.75
Average 135.11 12.57 29.92 24.75 24.63 5.30 18.11 15.29 15.07 8.32 9.18 6.81
The results on road-network data follow the same overall trends observed with Euclidean distances. The GNN achieves the lowest RMSE and MAPE on hold-out district-cost test sets, followed by the SNN, FIG, and BHHD estimators. In terms of operational efficiency, the districts produced by GNN-ILS remain, on average, about 7% to 9% more cost-effective than those obtained with alternative baselines.
Other comparisons with reference solutions. As mentioned earlier, using a full SAA approach that solves multiple TSPs to evaluate district costs at each step would be prohibitively slow. As an alternative, we implemented another simplified baseline that considers a single demand realization scenario. With this method, each district cost evaluation reduces to solving a deterministic TSP. In our experiments, this approach underperformed the GNN-ILS by 1.88% and 1.82% on average for Euclidean instances with central and off-center depots, respectively, and by 2.71% in the road-network cases, highlighting the value of leveraging stochastic information. Nevertheless, it performed favorably compared to ILS with the BHHD, FIG, or SNN estimators, likely because, unlike CA formulas relying on district statistics, the single-scenario TSP approach does not systematically steer the districting method towards certain suboptimal geometrical characteristics such as compactness. Its main drawback, however, is computational: even a single TSP solution per move evaluation in a districting method is considerably slower than evaluating a CA formula or performing neural network inference, permitting very few local search descents for districting optimization over a given time.
We also investigated replacing the heuristic TSP evaluations based on the Lin–Kernighan algorithm with machine-learning-based estimators such as the approach proposed by Varol, Özener, and Albey (2024), which predicts tour lengths from given point coordinates. However, under our experimental configuration with tours of around 100 customers, the fast parametrization of LKH consistently finds near-optimal solutions (optimal in 92% of the cases and within 0.1% of the exact solution on average) within 0.1 seconds, whereas the point-based estimator requires complex feature generation and typically yields larger TSP-estimation errors (around 1%) at a slower runtime of 0.6 to 1.2 seconds. Consequently, current point-based TSP estimators provide no computational or accuracy advantage over LKH. By contrast, the trained GNN takes roughly 8ms for inference on a given district, and takes demand distribution information as input instead of being tied to the inherent variability of a single demand scenario.
Finally, to broaden our comparison, we evaluated a constrained version of K-means (CK-means) using the algorithm of Levy-Kramer (2018). This method is extremely fast, requiring only a few seconds and no parameter calibration, but its solutions were, on average, 4.46% and 3.31% worse than those obtained with GNN-ILS in the Euclidean and road-network cases, respectively. Moreover, since the algorithm does not guarantee spatial connectivity, it frequently produced infeasible district configurations.
Generalization capabilities. Building on the architecture proposed in this work, a subsequent study by Ahmed et al. (2024) also extended the methodology into an end-to-end, decision-aware learning framework that integrates a combinatorial optimization layer within the GNN pipeline. In this formulation, the network parametrizes a surrogate problem based on a capacitated minimum spanning tree and is trained using a Fenchel-Young loss with perturbation-based gradient estimation following Dalle et al. (2022), allowing the model to optimize the downstream decision objective directly rather than minimizing district-cost prediction error. This end-to-end approach demonstrated strong generalization across cities, across varying demographic and geographical characteristics of BUs (e.g., population density, area, and compactness), and across problem scales. It also achieved improved computational efficiency by learning on small problem instances and extrapolating to much larger ones: a model trained on instances with about 30 BUs successfully producing high-quality districting solutions for metropolitan regions containing up to 2,000 units. Overall, while the architecture introduced in this paper constitutes a fundamental building block for efficient districting-and-routing methods, subsequent advances in learning have further extended its generalization capabilities.
Summary. Table 11 summarizes the performance, speed, and generalization properties of all the evaluated methods. The purpose of this study is not to present ILS-GNN as a universal tool to every districting-and-routing setting, but to highlight its robustness and the limitations of using CA formulas within optimization frameworks that tend to favor districts whose costs are most underestimated. The proposed approach provides a strong alternative whenever some preprocessing is acceptable and optimization must be repeated over multiple problem variants or partial solutions.
Table 11: Summary of the considered district-cost estimation and optimization methods
Method Learning Optimization Gap(%) Generalization to other cities
ILS-BHHD <<1s 20min 10.79 Retraining is optional – Fine-tuning regression weights for better performance requires district-cost examples
ILS-FIG <<1s 20min 11.03 Retraining is optional – Fine-tuning regression weights for better performance requires district-cost examples
ILS-SNN 1s 20min 10.69 Retraining using district-cost examples
ILS-GNN 2h 20min Retraining using district-cost examples, possible to generalize to different cities using end-to-end learning (Ahmed et al. 2024)
ILS-TSP 20min 1.83 No learning needed
CK-means a few seconds 4.46 No learning needed
Gap (%) relative to ILS-GNN. Euclidean Case. All districting solutions were evaluated on hold-out test scenarios.
 Single-scenario evaluation using LKH3 for the current fastest and most accurate estimations of TSP costs.
 Solutions might not respect district connectivity constraints.
@@ 7 Conclusions
This paper has investigated a novel approach for strategic districting and routing, relying on a GNN for district cost estimation within an ILS, aiming to find a good partition of a metropolitan area into districts. The GNN is trained through a supervised learning process, using pre-computed TSP costs for various districting configurations derived from the LKH heuristic. Our experimental analyses demonstrate that this approach generates efficient districts, leading to substantial economic gains of 10.12% on average (and exceeding 20% in some cases) compared to other commonly used methods for routing cost estimations. They also show that compactness is an insufficient proxy for efficiency and that other geometric characteristics (e.g., elongated district shapes towards the depot) play a key role in determining long-term operational costs.
This research opens many promising research avenues. A natural direction for future work is to analyze the applicability of GNN-based cost estimations to other stochastic problems that require time-intensive second-stage evaluations. One main advantage of the proposed GNN approach is its flexibility. It is driven by data (geographical features of the BUs and training examples) and does not require distributional information over customer requests. Consequently, it is fairly generic, and adapting it to other district-cost definitions (e.g., considering a different subproblem than a TSP, introducing additional constraints or network characteristics) boils down to replacing the approach used to generate labeled districts.
Another line of research involves further improving the learning methodology and its generalization capabilities. While the present work trains the GNN to minimize prediction error independently of the downstream optimization process, subsequent developments have shown that end-to-end, decision-aware training (using a surrogate optimization layer and a Fenchel–Young loss) can further enhance generalization across cities with different population densities, spatial structures, and district sizes (Ahmed et al. 2024).
Finally, a complementary perspective is to explore solution methods at the districting level that explicitly exploit finer properties of district costs. In many practical cases, adding an additional basic unit (BU) to a district leads to diminishing marginal increases in cost, suggesting an approximately submodular behavior of the district-cost function that could be leveraged algorithmically. However, submodularity violations might also arise when the inclusion of new BUs induces threshold effects, such as capacity or time-window constraints in the underlying VRP subproblems that require an additional vehicle, multi-depot settings where the optimal depot assignment changes, or spatial discontinuities introduced by bridges, tunnels, or other natural barriers. Investigating different optimization paradigms, approximation algorithms, or local-search heuristics that take advantage of specific structures to filter moves could provide valuable insights to improve solution methods, complementing the learning-based approaches examined in this study.